OpenStudy (anonymous):

what are the factors of 60a #hero

OpenStudy (anonymous):

please someone help

OpenStudy (anonymous):

what are the numbers that evenly divide into 60?

OpenStudy (anonymous):

$60 $30

OpenStudy (anonymous):

There are lots more. Start at 1 and keep going until you get a pair that repeats. 1 * ___ = 60 2 * ___ = 60 3* ___ = 60

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

you're just looking for numbers that multiply to give 60, like from your multiplication tables... For example: Find the factor of 24 1 * 24 = 24 2 * 12 = 24 3 * 8 = 24 4 * 6 = 24 5 doesn't go into 24 6 * 4 = 24 → 6 * 4 is the same thing as 4 * 6, so we can stop looking for factors. Therefore the factors of 24 are 1, 2, 3, 4, 6, 8, 12, and 24. Try it with 60

OpenStudy (anonymous):

ok